# Evaluating the Fitness of Ontologies for the Task of Question Generation

### v1

@@ 4 Response Analysis and Requirement Derivation
Observation: The first analysis examined whether the ontologies could generate all question types listed in Table 1. The Geography ontology was the largest considered (in terms of axioms), and consequently generated the highest number of questions (Table 4), although the majority of these were property-based. However, size alone does not determine its suitability for the task; whilst obviously influential, structural complexity is also equally (if not more critically) important in accommodating the diverse axiom requirements of different question types. For example, no annotation-based questions were generated for Geography as it lacked annotation properties that could match the pattern <x> <rdfs:comment> <string>. The Solar System ontology generated questions for each of the types, reflecting a comprehensive structure and completeness in satisfying the question pattern pre-requisites from Table 1. This contrasts with African Wildlife, which suffered from a lack of individuals, and only generated terminology-based and annotation-based questions. These observations emphasise that structural complexity, including the presence of all prerequisites (e.g. the patterns in Table 1) are essential for generating a comprehensive range of question types. Missing axioms or incomplete structures can significantly limit the ontology’s ability to support specific templates, hindering its effectiveness for question generation.
Requirement: An input ontology should assist in generating all desirable question types, regardless of the axiom prerequisites required for each question. This requirement can be evaluated through metrics that address two key questions:
@@
    “How many RDF pattern fragments align with the components of the ontology schema and metadata model?”
@@
Table 4: Number of questions generated for each question generation strategy.
Solar System Geography African Wildlife
Class membership 70 713 0
Property-based 120 2044 0
Terminology-based 9 1 25
Annotation-based 10 0 15
MCQs 8 0 0
Multi-entity 25 724 0
Total 242 3482 40
@@
Metric Pattern Coverage ($PC\in[0,1]$): The first question assesses the completeness of the ontology based on its utilisation of RDF pattern fragments (Table 1), described by the instantiation of concepts ($inst^{\mathcal{C}}$) and objects ($inst^{\mathcal{P}}$), the class hierarchy ($\mathcal{H}^{\mathcal{C}}$) and schema annotations ($\mathcal{A}$). The metric, defined as $PC=\mathcal{F}_{\text{used}}/\mathcal{F}_{\text{tot}}$ quantifies the ratio of pattern fragments used ($\mathcal{F}_{\text{used}}$) to the total available defined within the AQG patterns ($\mathcal{F}_{\text{tot}}$), indicating how effectively the ontology leverages its formal language. A high score suggests full pattern utilisation, while a lower score identifies missing elements that may hinder question generation (see Table 5, together with the scores for each ontology evaluated).
@@
Solar System achieved full coverage ($PC=1$), aligning with its ability to generate all question types (Table 4). Geography scored lower due to missing pattern types, particularly the absence of annotation properties (rdfs:comment), leading to $\mathcal{A}=\emptyset$. African Wildlife covered only half of the patterns, limiting its question generation to taxonomic and annotation types.
@@
Inheritance Richness $IR=\tfrac{1}{|\mathcal{C}|}\sum_{\mathcal{C}_{i}\in\mathcal{C}}\left|\mathcal{H}^{\mathcal{C}}(\mathcal{C}_{1},\mathcal{C}_{i})\right|$ 0.9 0.1 1.2
@@
Relationship Richness $\overline{RR}=\tfrac{1}{|\mathcal{C}|}\sum_{\mathcal{C}_{i}\in\mathcal{C}}RR_{C}$ 0.8 0.9 0
Average Connectivity $\overline{Cn}=\tfrac{1}{|\mathcal{C}|}\sum_{\mathcal{C}_{i}\in\mathcal{C}}Cn(\mathcal{C}_{i})$ 4.8 216 0
Sibling fan-outness $SF=\frac{|C^{\prime}_{\text{sib}}|}{|C|}$ 8 0 0
Average Depth $\overline{D}=\frac{1}{|\mathcal{C}|}\sum_{\mathcal{C}_{i}\in\mathcal{C}}CD(\mathcal{C}_{i})$ 2.3 1.1 2.3
Metric Class Richness ($CR\in[0,1]$): The second question (“What is the structural complexity level of each component in the ontology?”) extends the first by assessing structural complexity at both schema and data levels through two key metrics: Class Richness and Average Population (adopted from Tartir et al. [38, 39]). Class Richness ($CR=|\mathcal{C}^{\prime}|/|\mathcal{C}|$) measures the ratio of classes in the ontology ($\mathcal{C}$) that have been instantiated within the Knowledge Base ($\mathcal{C}^{\prime}$); which is essential for generating class-based questions (Table 5). A low $CR$ indicates a conceptually rich schema but insufficient KB instances for question generation.
Metric Average Population ($P\in\mathbb{R}$): measures the average number of instances per class, indicating whether classes are sufficiently populated ($P=|\mathcal{I}|/|\mathcal{C}|$). This directly affects class-based question generation and other instance-dependent templates, such as property-based strategies. Together with $CR$, it reveals the ontology population distribution, influencing both question diversity and volume.
For example, Geography (79 instances per class) generated 713 questions, whereas Solar System (7 instances per class) produced only 70.
Metric Inheritance Richness ($IR\in\mathbb{R}$): As terminology-based questions are solely dependent on the taxonomic relationships defined within the ontology, it is necessary to quantify the ratio between the total number of subclasses ($|\mathcal{C}_{1}|$) and the total number of classes $|\mathcal{C}|$ defined in the ontology schema. This metric determines the average number of subclasses per class ($IR={|\mathcal{C}|}^{-1}\sum_{\mathcal{C}_{i}\in\mathcal{C}}\left|\mathcal{H}^{\mathcal{C}}(\mathcal{C}_{1},\mathcal{C}_{i})\right|$).
African Wildlife achieved the highest score ($IR=1.2$) reflecting the higher number of terminology-based questions generated overall (Table 4).
Metric Relationship Diversity ($RD\in[0,1]$): To evaluate the ontology’s capability to generate property-based questions, two additional metrics are examined: Relationship Diversity and Relationship Richness. This question generation strategy relies on the use of properties $\mathcal{P}$ to connect objects $Ob=\mathcal{I}\cup L$ (i.e. instances $\mathcal{I}$ or literals $L$) thus targeting the instantiated properties $inst^{\mathcal{P}}$ in the A-Box, resulting in the formation of more detailed questions regarding concepts’ properties. Relationship Diversity reflects the balance between taxonomic and non-taxonomic relations, and is defined as $RD=|inst^{\mathcal{P}}|~{}/~{}(|\mathcal{H}^{\mathcal{C}}|+|inst^{\mathcal{P}}|)$. A high $RD$ indicates an emphasis on non-taxonomic relations, increasing property-based questions while reducing terminology-based ones.
Metric Relationship Richness ($\overline{RR}\in\mathbb{R}$): This quantifies the number of properties defined in the schema (T-box) for each class, that are utilised by instances in the KB (A-box) by finding the mean relationship richness for each class ($RR_{C}$). This is defined as the ratio of the number of properties $\mathcal{P}$ used by the instances $\mathcal{I}_{i}$ of class $\mathcal{C}_{i}$ (i.e. $|\mathcal{P}(\mathcal{I}_{i},Ob)|$) and the total number of properties defined for class $\mathcal{C}_{i}$ in the schema (i.e. $|\mathcal{P}(\mathcal{C}_{i},\mathcal{C}_{j})|$):
[…]
An ontology with a high $\overline{RR}$ (e.g. $\overline{RR}=0.8$ and $\overline{RR}=0.9$ for Solar System and Geography respectively) indicates a well-structured axiomatic foundation aligned with the prerequisites, highly suited for generating property-based questions.
4.2 RQ2: Supports the generation of multi-entity questions.
Observation: The Geography ontology generated the largest number of multi-entity questions (724 questions); e.g. “What is the highest point in Georgia with an elevation of 1,458 meters?”. The expert attributed this to the rich and diverse range of descriptions within that ontology relating to the breadth of properties used to model each concept; for example, the concept State in Geography included the properties statePopulation, stateArea and borders. Likewise, Solar System, though much smaller in size, generated 25 multi-entity questions, again due to having a broad set of properties. However, no multi-entity questions were generated for African Wildlife, as it lacked instances in its KB.
Requirement: The effectiveness of an ontology should consider support for the generation of the more advanced multi-entity questions. This type of question involves the use of more than one axiom, allowing for greater variation in question generation, thus testing learners’ knowledge of how concepts interact [37]. This variation is also important as it enables the creation of questions that go beyond simple, one-fact questions, offering a broader and deeper understanding of the domain (i.e. targetting the second level in Bloom’s taxonomy [8, 37]). Multi-entity questions tend to be more detailed and complex, and cover a wider breadth of knowledge within the domain. These questions are generated when an ontology contains rich concept descriptions, with multiple axioms that comprehensively define a concept by detailing its properties and associations [37].
Metric Average Connectivity ($\overline{Cn}\in\mathbb{R}$): The fact that multi-entity questions are formed by combining multiple facts regarding a concept poses the question “How well are instances of each class connected to objects?” Thus, a metric for this requirement should relate to Connectivity ($Cn$), which quantifies the out-degree (i.e. number of properties to other concepts) of a concept. This metric captures the centrality of classes within the ontology by quantifying how connected a class is to other objects ($Ob$), and is defined for properties $\mathcal{P}$ as:
[…]
Thus, at the ontology level, Average Connectivity ($\overline{Cn}$) measures the overall degree of interconnectivity across all classes in the ontology.
Geography scores highly ($\overline{Cn}=216$) and Solar System less so but with a positive score ($\overline{Cn}=5$), suggesting that both are well-suited for generating multi-entity questions, which typically involve 2–3 connections per question.
@@
Requirement: Multiple Choice Questions (MCQs) are the most common question types generated as they offer a more complex structure that requires the incorporation of several axioms with special characteristics. The structure of an ontology plays a key role in this process, particularly in the generation of plausible distractors for MCQs. These distractors are crafted to resemble the correct answer by leveraging semantic relationships within the ontology. Specifically, sibling classes (i.e. those that share a common immediate parent) are ideal candidates for distractors due to their inherent semantic similarity [13, 30]. In order to generate plausible distractors, an ontology should contain a hierarchical structure where populated sibling classes are well-represented, leading to the question “Does the ontology contain populated classes with a sufficient number of sibling classes?”.
Metric Sibling Fan-Outness ($SF\in\mathbb{R}$): This can be measured by calculating the ratio of the absolute populated sibling cardinality ($\mathcal{C}^{\prime}_{\text{sib}}$) to the total number of populated classes ($\mathcal{C}^{\prime}$) in the ontology; i.e. $SF=|C^{\prime}_{\text{sib}}|/|C|$. A high $SF$ score indicates that the ontology has a horizontally branching structure with multiple sibling classes under a few superclasses. This results in a broad hierarchy where categories are general rather highly differentiated. Such a structure is particularly effective for generating MCQ distractors with close semantic similarity.
Solar System was the only ontology to exhibit this pattern and successfully generated MCQs ($SF=8$), whereas as Geography lacked classes that shared a common parent, no MCQs were generated. Likewise, as African Wildlife lacked any instances in its KB, no MCQs were generated ($SF=0$).
@@
Metric Average Depth ($\overline{D}\in\mathbb{R}$): This requirement can be assessed by considering whether the ontology includes concepts from varying depths of the hierarchical structure. To evaluate this, $\overline{D}={|\mathcal{C}|}^{-1}\sum_{\mathcal{C}_{i}\in\mathcal{C}}CD(\mathcal{C}_{i})$ calculates the average depth $CD(\mathcal{C}_{i})$ of each concept $\mathcal{C}_{i}\in\mathcal{C}$, defined as the number of edges from $C_{i}$ to the root in hierarchy $H^{c}$. This metric (similar to that defined by Gangemi et al. [16]) ensures question generation spans general to specific concepts, highlighting that hierarchical richness, rather than size alone, drives cognitive diversity.
@@
Both Solar System and African Wildlife (each score $\overline{D}$ = 2.3) exhibited rich hierarchies, generating questions at varying conceptual levels. For example, Solar System included broad concepts like “Planet” while deeper levels introduced distinctions such as “Terrestrial Planet” and “Gas Giant”.

### v2

@@ 4 Response Analysis and Requirement Derivation
Observation: The first analysis examined whether the ontologies could generate all of the question types listed in Table 1. The Geography ontology was the largest considered (in terms of axioms), and consequently generated the highest number of questions (Table 3), although the majority of these were property-based. However, size alone does not determine its suitability for the task; whilst obviously influential, structural complexity is also equally (if not more critically) important in accommodating the diverse axiom requirements of different question types. For example, no annotation-based questions were generated for Geography as it lacked annotation properties that could match the pattern <x> <rdfs:comment> <string>. The Solar System ontology generated questions for each of the types, reflecting a comprehensive structure and completeness in satisfying the question pattern pre-requisites from Table 1. This contrasts with African Wildlife, which suffered from a lack of individuals, and only generated terminology-based and annotation-based questions. These observations emphasise that structural complexity, including the presence of all prerequisites (e.g. the patterns in Table 1) are essential for generating a comprehensive range of question types. Missing axioms or incomplete structures can significantly limit the ontology’s ability to support specific templates, hindering its effectiveness for question generation.
Requirement: An input ontology should assist in generating all desirable question types, regardless of the axiom prerequisites required for each question. This requirement can be evaluated through several metrics that address two key questions:
@@
    “How many pattern fragments align with the components of the ontology schema and metadata model?”
@@
Metric: Pattern Coverage ($PC\in[0,1]$). The first question assesses the completeness of the ontology based on its utilisation of RDF pattern fragments (Table 1), described by the instantiation of concepts ($inst^{\mathcal{C}}$) and objects ($inst^{\mathcal{P}}$), the class hierarchy ($\mathcal{H}^{\mathcal{C}}$) and schema annotations ($\mathcal{A}$). The metric, defined as $PC=\mathcal{F}_{\text{used}}/\mathcal{F}_{\text{tot}}$ quantifies the ratio of pattern fragments used ($\mathcal{F}_{\text{used}}$) to the total available defined within the AQG patterns ($\mathcal{F}_{\text{tot}}$), indicating how effectively the ontology leverages its formal language. A high score suggests full pattern utilisation, while a lower score identifies missing elements that may hinder question generation (see Table 5, together with the scores for each ontology evaluated).
@@
Solar System achieved full coverage ($PC=1$), aligning with its ability to generate all question types (Table 3). Geography scored lower due to missing pattern types, particularly the absence of annotation properties (rdfs:comment), leading to $\mathcal{A}=\varnothing$. African Wildlife covered only half of the patterns, limiting its question generation to taxonomic and annotation types.
Metric: Class Richness ($CR\in[0,1]$). The second question (“What is the structural complexity level of each component in the ontology?”) extends the first by assessing structural complexity at both schema and data levels through two key metrics: Class Richness and Average Population (adopted from Tartir et al. [38, 39]). Class Richness ($CR=|\mathcal{C}^{\prime}|/|\mathcal{C}|$) measures the ratio of classes in the ontology ($\mathcal{C}$) to those that have been instantiated within the Knowledge Base ($\mathcal{C}^{\prime}$); which is essential for generating class-based questions (Table 5). A low $CR$ indicates insufficient instantiation of classes within the KB, limiting its suitability for question generation.
Metric: Average Population ($P\in\mathbb{R}$). This metric measures the average number of instances per class, indicating whether classes are sufficiently populated ($P=|\mathcal{I}|/|\mathcal{C}|$). This directly affects class-based question generation and other instance-dependent templates (e.g. property-based strategies). Together with $CR$, it reveals the ontology population distribution, influencing both question diversity and volume.
For example, Geography (79 instances per class) generated 713 questions, yet Solar System (7 instances per class) produced only 70.
Metric: Inheritance Richness ($IR\in\mathbb{R}$). As terminology-based questions are solely dependent on the taxonomic relationships defined within the ontology, it is necessary to quantify the ratio between the total number of subclasses ($|\mathcal{C}_{1}|$) and the total number of classes $|\mathcal{C}|$ defined in the ontology schema. This metric determines the average number of subclasses per class ($IR={|\mathcal{C}|}^{-1}\sum_{c\in\mathcal{C}}\left|\mathcal{H}^{\mathcal{C}}(c_{1},c_{2})\right|$).
Solar System achieved the highest score ($IR=0.9$), indicating a denser use of subclassing despite its smaller size, while African Wildlife scored slightly lower ($IR=0.8$) due to a larger number of top-level classes.
Metric: Relationship Diversity ($RD\in[0,1]$). To evaluate the ontology’s capability to generate property-based questions, two additional metrics are examined: Relationship Diversity and Relationship Richness. This question generation strategy relies on the use of properties $\mathcal{P}$ to connect objects $Ob=\mathcal{I}\cup L$ (i.e. instances $\mathcal{I}$ or literals $L$) thus targeting the instantiated properties $inst^{\mathcal{P}}$ in the knowledge base, resulting in the formation of more detailed questions regarding concepts’ properties. Relationship Diversity reflects the balance between taxonomic and non-taxonomic relations, and is defined as $RD=|inst^{\mathcal{P}}|~/~(|\mathcal{H}^{\mathcal{C}}|+|inst^{\mathcal{P}}|)$. A high $RD$ indicates an emphasis on non-taxonomic relations, increasing property-based questions while reducing terminology-based ones.
Metric: Average Relationship Richness ($\overline{RR}\in\mathbb{R}$). This quantifies the number of properties defined in the schema for each class and that are utilised by instances in the KB, by first computing a relationship richness value for each class ($RR_{c_{i}}$), and then taking the mean across all classes. This is defined as the ratio of the number of properties used by the instances of class $c_{i}$ (i.e. $|p(i,o)|$) and the total number of properties defined for class $c_{i}$ in the schema (i.e. $|p(c_{i},c_{j})|$):
$\overline{RR}=\frac{1}{|\mathcal{C}|}\sum_{c_{i}\in\mathcal{C}}RR_{\mathcal{C}},\text{~~~~where~}RR_{c_{i}}=\frac{|p(i,o)|,i\in c_{i}(\mathcal{I})}{\left|p(c_{i},c_{j})\right|}$
An ontology with a high $\overline{RR}$ (e.g. $\overline{RR}=0.5$ and $\overline{RR}=0.9$ for Solar System and Geography respectively) indicates a well-structured axiomatic foundation aligned with the prerequisites, highly suited for generating property-based questions.
4.2 RQ2: Supports the generation of multi-entity questions.
Observation: The Geography ontology generated the largest number of multi-entity questions (724 questions); e.g. “What is the highest point in Georgia with an elevation of 1,458 meters?”. The expert attributed this to the rich and diverse range of descriptions within that ontology relating to the breadth of properties used to model each concept; for example, the concept State in Geography included the properties statePopulation, stateArea and borders. Likewise, Solar System, though much smaller in size, generated 17 multi-entity questions, again due to having a broad set of properties. However, no multi-entity questions were generated for African Wildlife, as it lacked instances in its KB.
Requirement: The effectiveness of an ontology should consider support for the generation of the more advanced multi-entity questions. This type of question involves the use of more than one axiom, allowing for greater variation in question generation, thus testing learners’ knowledge of how concepts interact [37]. This variation is also important as it enables the creation of questions that go beyond simple, one-fact questions, offering a broader and deeper understanding of the domain (i.e. targeting the second level in Bloom’s taxonomy [8, 37]). Multi-entity questions tend to be more detailed and complex, and cover a wider breadth of knowledge within the domain. These questions are generated when an ontology contains rich entity descriptions, with multiple axioms that comprehensively define an entity by detailing its properties and associations [37].
Metric: Average Connectivity ($\overline{Cn}\in\mathbb{R}$). The fact that multi-entity questions are formed by combining multiple facts regarding an entity poses the question “How well are instances connected to other objects?” The corresponding metric should therefore capture the notion of Connectivity ($Cn$), which quantifies the out-degree of an instance (i.e the number of property assertions it makes). This reflects how richly described instances are within the ontology, and is defined as:
$\overline{Cn}=\frac{1}{|\mathcal{I}|}\sum_{i\in I}|p(i,o)|$
where $\mathcal{I}$ is the set of all instances in the ontology, and $p(i,o)$ denotes the set of outgoing property assertions (both object and datatype) made by instance $\mathcal{I}_{i}$. Thus, Average Connectivity captures the overall degree of factual richness at the instance level.
Geography scores highly ($\overline{Cn}=2.8$) and Solar System less so but with a positive score ($\overline{Cn}=1.7$), suggesting that both are well-suited for generating multi-entity questions, which typically involve 2–3 connections per question.
@@
Inheritance Richness $IR=\tfrac{1}{|\mathcal{C}|}\sum_{c\in\mathcal{C}}\left|\mathcal{H}^{\mathcal{C}}(c_{1},c_{2})\right|$ 0.9 0.1 0.8
@@
Av. Relationship Richness $\overline{RR}=\tfrac{1}{|\mathcal{C}|}\sum_{\mathcal{C}_{i}\in\mathcal{C}}RR_{C}$ 0.5 0.9 0
Average Connectivity $\overline{Cn}=\frac{1}{|\mathcal{I}|}\sum_{i\in I}|p(i,o)|)$ 1.7 2.8 0
Sibling fan-outness $SF=\frac{|C^{\prime}_{\text{sib}}|}{|C|}$ 0.8 0 0
Average Depth $\overline{D}=\frac{1}{|\mathcal{C}|}\sum_{\mathcal{C}_{i}\in\mathcal{C}}CD(\mathcal{C}_{i})$ 2.3 1.1 1.9
@@
Requirement: MCQs are the most common question types generated as they offer a more complex structure that requires the incorporation of several axioms with special characteristics. The structure of an ontology plays a key role in this process, particularly in the generation of plausible distractors for MCQs. These distractors are crafted to resemble the correct answer by leveraging semantic relationships within the ontology. Specifically, sibling classes (i.e. those that share a common immediate parent) are ideal candidates for distractors due to their inherent semantic similarity [13, 30]. In order to generate plausible distractors, an ontology should contain a hierarchical structure where populated sibling classes are well-represented, leading to the question “Does the ontology contain populated classes with a sufficient number of sibling classes?”.
Metric: Sibling Fan-Outness ($SF\in\mathbb{R}$). This is measured by calculating the ratio of the absolute populated sibling cardinality ($\mathcal{C}^{\prime}_{\text{sib}}$) to the total number of populated classes ($\mathcal{C}^{\prime}$) in the ontology; i.e. $SF=|C^{\prime}_{\text{sib}}|/|\mathcal{C}^{\prime}|$. A high $SF$ score indicates that the ontology has a horizontally branching structure with multiple sibling classes under a few superclasses. This results in a broad hierarchy where categories are general rather highly differentiated. Such a structure is particularly effective for generating MCQ distractors with close semantic similarity.
Solar System was the only ontology to exhibit this pattern and successfully generated MCQs ($SF=0.8$), whereas as Geography lacked classes that shared a common parent, no MCQs were generated. Likewise, as African Wildlife lacked any instances in its KB, it also generated no MCQs ($SF=0$).
@@
Metric: Average Depth ($\overline{D}\in\mathbb{R}$). This requirement can be assessed by considering whether the ontology includes concepts from varying depths of the hierarchical structure. To evaluate this, $\overline{D}={|\mathcal{C}|}^{-1}\sum_{c_{i}\in\mathcal{C}}CD(c_{i})$ is used to where $CD(c_{i})$ is the depth of concept $c_{i}$ in the hierarchy $H^{c}$ (defined as the edges from $c_{i}$ to the root). This metric (similar to that defined by Gangemi et al. [16]) yields a single average depth across all concepts to ensure question generation spans general to specific concepts, highlighting that hierarchical richness, rather than size alone, drives cognitive diversity.
@@
Both Solar System ($\overline{D}=2.3$) and African Wildlife ($\overline{D}=1.9$) exhibited rich hierarchies, generating questions at varying conceptual levels. For example, Solar System included broad concepts like “Planet” while deeper levels introduced distinctions such as “Terrestrial Planet” and “Gas Giant”.
5 Empirical Validation of the Proposed Metrics
To validate the metrics within the context of AQG, four ontologies that had previously been used in ontology-based question generation studies were selected (Music [12], Job [43], Restaurant [43] and People & Pets [4]) and evaluated both in terms of their profile (using the metrics summarised in Table 5), and the reported characteristics of the questions generated by each study. In each case, both the ontology and their associated knowledge bases were analysed, and the values for each metric were calculated (Table 6). The radar diagram (Figure 1) provides a comparative analysis of the normalised profiles (as defined by the different metrics) for each ontology. Higher values indicate more developed features, such as dense class instantiation, diverse relationships, or deeper class hierarchies. By analysing each of these profiles, we demonstrate that the metric scores (and hence the metric definitions themselves) align with and complement the prior findings of other independent studies, thus validating the relevance of the proposed metrics for ontology-based AQG approaches.
Cubric and Tosic [12] utilised the Music Ontology for question generation and reported notable limitations in producing questions that relied on Property-based and Class-based strategies (particularly MCQs). These generation strategies rely heavily on the availability of richly instantiated data, specifically well-defined class-instance relationships (e.g. <ex:TheBeatles> <rdf:type> <mo:MusicArtist>) and instance-property assertions (e.g. <ex:TheBeatles> <mo:genre> <ex:Rock>). However, such relationships are largely absent from the ontology itself, which primarily encodes schematic structures rather than populated instance data. This under-representation is directly reflected in the different values for the evaluation metrics in Table 6.
Music resulted in low scores across several key evaluation metrics. Notably, its Class Richness is limited, with only 3% of classes populated with instances. Moreover, the Average Population per class is extremely low, at just 0.02 instances per class. Compounding this, the Average Relationship Richness metric indicates that none of the properties defined at the schema level are utilised to describe instances at the data level. Collectively, these metrics reflect a low degree of instantiation and limited interconnection among instances within the ontology. This supports Cubric and Tosic’s observation that their AQG approach performed poorly on Music for Class-based and Property-based question generation, largely due to the absence of meaningful associations between classes and instances [12].
The results demonstrate that Music is primarily schematic (as further indicated by its relatively high Inheritance Richness, Relationship Diversity and Average Depth scores), encoding structural knowledge about musical concepts (e.g., Genre, Instruments and Lyrics) rather than detailed instance-level data. Consequently, it is better suited to supporting Terminology-based questions, which depend predominantly on class hierarchies and labels, and is less appropriate for generating Property-based, Class-based or Multi-entity questions that require a richer population of instances. Thus, the metric results align closely with the ontology’s reported performance in prior AQG experiments.
Vinu and Kumar [43] utilised the Job and Restaurant ontologies for Class-based and Property-based question generation. They reported a high number of relevant triples (triples retrieved from the ontologies that satisfy question requirements); for example, over 288,000 triples were retrieved from the Restaurant ontology using a single question pattern. These triple counts serve as an indicator of the ontologies’ potential to support large-scale question generation and reflect their extensive instance populations.
Figure 1: Normalised results for the proposed ontology evaluation metrics applied to the Job, Restaurant, Music, and People & Pets ontologies
This observation is supported by the results of the evaluation metrics in Table 6, which show that both ontologies exhibit high Class Richness and Average Population scores. The high Class Richness scores for Job ($CR=0.8$), and for Restaurant ( $CR=1$) indicate that almost all classes are populated with instances. Both Restaurant and Job are also sufficiently populated; however, the metrics for Restaurant suggest a better utilisation of defined properties at the schema level, with a Average Relationship Richness of $\overline{RR}=0.5$ compared to Job ($\overline{RR}=0.1$), which indicates a higher level of instance interconnection. The structural analysis also reveals that both schemata are relatively flat. In particular, the Sibling Fan-Outness metric is low, suggesting that although the classes are well populated, the ontologies lack a broad class hierarchy with a sufficient number of sibling classes. This limitation has practical implications. Whilst Job and Restaurant are well suited to generating question stems for Class-based and Property-based questions, they are less effective for generating MCQs with semantically plausible distractors, a process that typically relies on the presence of sibling classes (Section 4.3). Notably, Vinu and Kumar’s study [43] did not attempt full MCQ generation with distractors from these ontologies, focusing instead on stem generation, which is consistent with the structural characteristics revealed by this metric-based analysis.
The People & Pets ontology was employed by Alsubait et al. [4] to generate analogy-style, Terminology-based MCQs that rely on structural patterns such as class–subclass and sibling relationships. As illustrated in Table 6, the evaluation metrics align with this intended use, but also explain the relatively limited output (only 15 questions were generated from this ontology) [4]. The Inheritance Richness score ($IR=0.4$) and Average Depth ($\overline{D}=1.5$) suggest that the ontology possesses only a moderate hierarchical structure, which probably constrained the diversity of eligible concept pairs, thus contributing to the small number of generated analogies.
More critically, the ontology exhibits a high Sibling Fan-Outness score ($SF=0.5$), indicating that several classes are positioned under shared superclasses with sufficient horizontal breadth. This structure is particularly valuable for generating MCQs with semantically plausible distractors, as distractors in analogy questions often draw on sibling concepts to maintain a coherent semantic frame. This structural trait likely played a central role in enabling the generation of MCQs, even if the overall quantity was low. The ontology’s low Class Richness ($CR=0.2$) and Average Population ($P=0.3$) further confirm that it is schema-focused and sparsely populated with instances. Additionally, the Relationship Diversity and Average Relationship Richness scores ($RD=0.3$ and $\overline{RR}=0.01$, respectively) are low, reinforcing the notion that the ontology supports limited relational inference beyond subclass structures. Thus, the People & Pets ontology demonstrates structural properties suitable for Terminology-based MCQ generation. While the relatively shallow hierarchy limited the number of analogies that could be extracted, the strong presence of populated sibling structures made it especially effective for generating MCQs with meaningful distractors [4].
Table 6: Comparative evaluation of the validation ontologies across the proposed evaluation metrics.
Evaluation Metric Music Restaurant Job People & Pets
Pattern Coverage ($PC$) 1 0.8 1 1
Class Richness ($CR$) 0.03 1 0.8 0.2
Average Population ($P$) 0.2 2436 517 0.3
Inheritance Richness ($IR$) 0.9 0 0.1 0.4
Relationship Diversity ($RD$) 0 1 1 0.3
Av. Relationship Richness ($\overline{RR}$) 0 0.5 0.1 0.01
Average Connectivity ($\overline{Cn}$) 0 5.9 6 0.5
Sibling Fan-Outness ($SF$) 0 0 0 0.5
Average Depth ($\overline{D}$) 2.2 1 1.1 1.5